The cost of admission to the carnival was $50 for a group of 9 children and 2 adults. The admission was $61 for another group of 10 children and 3 adults. What was the admission for one child? For one adult?
Show work please

Answer by checkley79(3341)

9C+2A=50 MULTIPLY THIS EQUATION BY 3
10C+3A=61 MULTIPLY THIS EQUATUION BY -2 & ADD.
27C+6A=150
-20C-6A=-122
--------------
7C=28
C=28/7
C=$4 PER CHILD.
9*4+2A=50
36+2A=50
2A=50-36
2A=14
A=14/2
A=$7 FOR ADULT.
PROOF:
10*4=3*7=61
40+21=61
61=61